Find the next number in the sequence
1,15,169,1695 ?

Found 2 solutions by rothauserc, greenestamps:
Answer by rothauserc(4718)

1,15,169,1695, ?
:
The nth element of this sequence is 8^n - 7^n, where n varies 1, 2, 3, 4, 5, 6, ...
:
a(5) = 8^5 - 7^5 = 32768 - 16807 = 15961
:

Answer by greenestamps(13200)



I have no idea how the other tutor came up with a formula for the n-th term in your sequence....

However, in fact this is an invalid question, because you could put ANY number next and it would be a valid sequence.

The fact that there is a simple formula that produces the first four terms of the sequence makes his answer of 15961 a PROBABLE answer to the question; but it can't be said that it is THE answer.

For ANY question like this, where you are only given some numbers in a sequence, without any information on what kind of sequence it is, it is IMPOSSIBLE to know the "RIGHT" answer.